OpenStudy (anonymous):

if t/5 =7/r then r/7 =??

OpenStudy (anonymous):

5/t

OpenStudy (anonymous):

mind helping on another??

OpenStudy (skullpatrol):

Did you learn anything from the first one?

OpenStudy (anonymous):

i kinda get it i was going to ask him how he got it too lol

OpenStudy (anonymous):

if r/7 is flipped then you would do the same with the other equation because they are equal

OpenStudy (anonymous):

ok but how would i do it for x/7 =y/5 then 5x=? if their not just flipped would it be 7x ??

OpenStudy (anonymous):

it would be 7y because they flipped to multiplication and they switched variables

OpenStudy (anonymous):

ohh ok i get it what about when they ask you to solve the like 6/r = 2/3 would i just do it like solving a fraction??

OpenStudy (anonymous):

yes you would solve for r

OpenStudy (anonymous):

.so id cross multiply

OpenStudy (anonymous):

i think i got it is it 9

OpenStudy (anonymous):

yes it is :)

OpenStudy (anonymous):

thanks! i get it now (:

OpenStudy (anonymous):

your welcome

OpenStudy (skullpatrol):